iosangel:

There is 0.5 L of air in a rigid, sealed container at a pressure of 203 kPa and a temperature of 203 K. The container is heated to 273 K. What is the final pressure of the air?

loveboy:

Since, the container is rigid, there is no inflow as well as outflow of air. Therefore, Pressure and Temperature remains constant. Using Gay-Lussac's Law: The Pressure Temperature Law, \[\frac{P1}{T1}=\frac{P2}{T2}\] Plug the values into the equation, \[\frac{203k}{203k}=\frac{P2}{273k}\] \[P2=273k\] Thus, final pressure of air would be 273kPa.

loveboy:

:D